OpenStudy (anonymous):

peyton manning completed 38 passes for 418 yards. at this rate how many yards would be gained for 22 passes?

OpenStudy (anonymous):

@amistre64 not good with word problems :l

OpenStudy (amistre64):

we are given a ratio of passes to yards: 38: 418 we are asked to modify this by some value, x, such that 38x = 22 for the ratio to remain the same we modify both sides as: 38x:418x find x to determine the value of 418x

OpenStudy (anonymous):

11?

OpenStudy (amistre64):

x = 22/38 = 11/19

OpenStudy (anonymous):

so its 11/19?

OpenStudy (amistre64):

we are given a ratio of passes to yards: 38: 418 we are asked to modify this by some value, x, such that 38x = 22 for the ratio to remain the same we modify both sides as: 38x:418x find x to determine the value of 418x ..... we just determined that x=11/19, so use that to determine the rest

OpenStudy (anonymous):

idk what to do with 11/19 ... ._.